Selected information from the Blake Corporation accounting records for June follows:

Materials Inventory
BB (6/1) 94,000
466,000 421,000
Work-In-Process Inventory
Labor 585,000
EB(6/30) 590,000
Finished Goods Inventory
BB (6/1) 296,000
841,000 838,000
Cost of Goods Sold
29,000
Manufacturing Overhead Control
283,000
283,000
Applied Manufacturing Overhead
312,000
283,000
29,000

Additional information for June follows:

  • Labor wage rate was $30 per hour.
  • During the month, sales revenue was $1,039,000, and selling and administrative costs were $241,000.
  • This company has no indirect materials or supplies.
  • The company applies manufacturing overhead on the basis of direct labor-hours.  

Required:

a. What was the cost of direct materials purchased in June?

b. What was the over- or underapplied manufacturing overhead for June?

c. What was the manufacturing overhead application rate in June?

d. What was the cost of products completed during June?

e. What was the balance of the Work-in-Process Inventory account at the beginning of June?

f. What was the operating profit (or loss) for June? (Negative amounts should be indicated by a minus sign.)
Answer #1

a) Cost of direct material purchased = 466000

b) Over applied overhead = 312000-283000 = 29000

c) Manufacturing overhead rate = 312000/19500 = 16 per labor hour

d) Cost products completed = 838000

e) ending WIP = 590000+838000-421000-312000-585000 = 110000

f) Operating profit = 1039000-(838000-29000)-241000 = -11000
Answer #2

B) 585000/30= 19500